SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (April 11, 2006)

Attendees: Kris Chaisanguanthum, Al Eisner, Fang Fang, Arik Kreisel, Selina Li, Rahmat Rahmat

Operations

  • BaBar wasn't taking much data recently because of water leak. We should be back taking colliding beam data again sometime this evening.
  • The DCZ failed to configure for a few times over the past week during data taking. See the DCT/DCZ section below for more details.

EMT

  • There was a special meeting to discuss the EMT deadtime last Friday afternoon. Chris O'Grady showed us that the fiber transfer takes about 60 microseconds while the FEX takes 54 microseconds which are both small compared to the simulation model predicts for 4kHz. It was suspected that the deadtime might be due to the ODF side. Chris and Matt Weaver are investigating this possibility.
  • In run 63581 taken on April 5th, there was 100% deadtime in EMT. The run was stopped and things were back to normal after a reconfigure.

DCT/DCZ

  • For the past week, there were 3 incidents which required power cycling of the TSF crates due to a missing board and thus DCZ failed to configure. In all incidents, the user reset mode was on and DCM wasn't locked. One incident was due to TSF-Y2, and the other times were due to TSF-X6 and TSF-X12. It looked like it could happen to any boards and any TSF crates. Marc suspected that the clock coming into the FPGA has too much noise/jitter for the DCMs to lock to. He is currently trying to find out why the clock isn't locking. Marc suggested that we can either switch back to the previous version of the firmware without the user reset function or change the online code to increase the times we hold the reset signal and also the time we allow for the clock to lock, before toggling the reset line again and repeating the loop. We decided to first try putting in a longer delay after the user reset in the ROM code.
  • There were some TSF damages in run 63680 (April 5th) and out-of-order error messages were seen in the TSF-X ROM logfile. Fang is looking at the headers for the TSF to find clues to the cause of the damages.
  • The old DCT system has been out of data taking since run 63595 (April 5th). The three old DCT crates were switched off on April 7th.

Data Quality and Monitoring

  • Rahmat showed us some lumi strip charts that he is working on. It looked like the error bars in the plots are not accurate yet. Al asked about the possibility of having these strip charts for data points per shift rather than per run. Rahmat will ask the DQG about tools for doing this.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.